# Composite Indicator-Guided Infilling Sampling for Expensive Multi-Objective Optimization

### v1

@@ IV-H Computational Complexity Analysis
In addition, we compared the running time of CI-EMO with five SAEAs on DTLZ2 problems with 2 and 3 objectives, as shown in Figure 6. Due to CI-EMO samples only one candidate in each iteration, its running time is relatively high. But the running time of CI-EMO is less than EMMOEA and R2DEGO on 3- objective DTLZ2 problem, where R2DEGO samples 5 samples in each iteration. Overall, the running time of CI-EMO is comparable to other SAEAs and is reasonable for expensive optimization scenarios since expensive fitness evaluation is time-consuming, for example, once computational fluid dynamics simulations could take hours [31].
@@
TABLE VII: HV STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FIVE SAEAS ALGORITHMS ON THREE REAL-WORLD PROBLEMS
@@
IV-I Comparison with state-of-the-art SAEAs on real-world problems
In this section, we empirically evaluate our proposed method based on three real-world MOPs: gear train design problem [46], car side impact design problem [47], and two bar plane truss [48], the detailed description of problems can be seen in the references due to space constraints. Their objective and decision variable numbers are shown in Table VII. We compared our approach with five advanced SAEAs. For each algorithm, the number of initial samples $\mathcal{N}_{0}$ is set to 100, other experimental setup was consistent with our previous study, and all experiments are conducted over 21 independent runs to ensure statistical reliability. To summarize the performance, we recorded the hypervolume values obtained for each method, which are reported in Table VII.
@@ V CONCLUSION
TABLE S-II: IGD+ RESULTS (MEAN AND STANDARD DEVIATION) OBTAINED BY THE PROPOSED CI-EMO AND THE VARIANT WITH SAME WEIGHTS

### v2

@@ 4.8 Computational Complexity Analysis
In addition, we compared the running time of CI-EMO with five SAEAs on DTLZ2 problems with 2 and 3 objectives, as shown in Figure 6. Due to CI-EMO samples only one candidate in each iteration, its running time is relatively high. But the running time of CI-EMO is less than EMMOEA and R2DEGO on 3- objective DTLZ2 problem, where R2DEGO samples 5 samples in each iteration. Overall, the running time of CI-EMO is comparable to other SAEAs and is reasonable for expensive optimization scenarios since expensive fitness evaluation is time-consuming, for example, once computational fluid dynamics simulations could take hours (Song et al., 2021).
@@
Table 7: HV STATISTIC RESULTS (MEAN AND STANDARD DEVIATION) FOR CI-EMO AND OTHER FIVE SAEAS ALGORITHMS ON THREE REAL-WORLD PROBLEMS
@@
4.9 Comparison with state-of-the-art SAEAs on real-world problems
In this section, we empirically evaluate our proposed method based on three real-world MOPs: gear train design problem (Ray and Liew, 2002), car side impact design problem (Jain and Deb, 2013), and two bar plane truss Coello Coello and Pulido (2005), the detailed description of problems can be seen in the references due to space constraints. Their objective and decision variable numbers are shown in Table 7. We compared our approach with five advanced SAEAs. For each algorithm, the number of initial samples $\mathcal{N}_{0}$ is set to 100, other experimental setup was consistent with our previous study, and all experiments are conducted over 21 independent runs to ensure statistical reliability. To summarize the performance, we recorded the hypervolume values obtained for each method, which are reported in Table 7.
@@ 5 CONCLUSION
Acknowledgment
This work was partly supported by the National Natural Science Foundation of China under Grant Nos. 62403442 and 42327803, the Postdoctoral Fellowship Program of CPSF under Grant No. GZC20241600, the "CUG Scholar" Scientific Research Funds at China University of Geosciences (Wuhan) (Project No. 2024011), the Fundamental Research Founds for National University, China University of Geosciences (Wuhan), as well as support from the Hubei Province Postdoctoral Talent Introduction Program (Project No. 2024HBBHJD096).
@@ Acknowledgment
Table S-II: HV RESULTS (MEAN AND STANDARD DEVIATION) OBTAINED BY THE PROPOSED CI-EMO AND OTHER FIVE SAEAS ALGORITHMS ON MaF BENCHMARK PROBLEMS
Problem M D KRVEA KTA2 EMMOEA DirHVEGO R2DEGO CI-EMO
MaF1 3 10 1.9027e-1 (3.99e-3) - 2.1218e-1 (6.78e-3) - 2.1419e-1 (2.26e-3) - 2.0125e-1 (2.85e-3) - 1.9899e-1 (3.23e-3) - 2.2107e-1 (1.82e-3)
5 10 8.5585e-3 (8.71e-4) - 7.9485e-3 (7.68e-4) - 1.0061e-2 (5.99e-4) $\approx$ 6.4275e-3 (2.50e-4) - 8.0139e-3 (2.94e-4) - 1.0316e-2 (3.64e-4)
10 10 1.0202e-7 (2.06e-8) - 2.1497e-7 (9.91e-8) - 5.1605e-7 (7.58e-8) - 1.0458e-7 (2.10e-8) - 1.1104e-7 (1.97e-7) - 7.2391e-7 (1.27e-7)
MaF2 3 10 2.3310e-1 (4.62e-3) - 2.4280e-1 (1.30e-3) - 2.3329e-1 (3.76e-3) - 2.4737e-1 (3.67e-4) + 2.4169e-1 (1.20e-3) - 2.4492e-1 (1.27e-3)
5 10 1.8142e-1 (3.24e-3) - 1.9020e-1 (2.02e-3) - 1.8425e-1 (3.26e-3) - 1.9658e-1 (1.28e-3) $\approx$ 1.9235e-1 (2.23e-3) - 1.9576e-1 (2.68e-3)
10 10 2.1822e-1 (3.39e-3) - 2.0633e-1 (2.29e-3) - 2.1672e-1 (4.26e-3) - 2.1638e-1 (2.38e-3) - 2.2898e-1 (1.58e-3) + 2.2122e-1 (2.25e-3)
MaF3 3 10 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0)
5 10 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0)
10 10 5.0320e-2 (1.11e-1) $\approx$ 6.1107e-1 (3.42e-1) + 1.3805e-1 (1.95e-1) $\approx$ 2.4563e-1 (2.49e-1) + 3.3580e-1 (3.05e-1) + 9.9335e-2 (2.48e-1)
MaF4 3 10 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0)
5 10 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0) $\approx$ 0.0000e+0 (0.00e+0)
10 10 1.1402e-7 (1.83e-7) - 1.7436e-6 (2.39e-6) $\approx$ 1.1127e-5 (1.66e-5) $\approx$ 2.3326e-7 (4.86e-7) - 6.9921e-7 (1.19e-6) - 3.5638e-6 (4.31e-6)
MaF5 3 10 5.8546e-2 (5.51e-2) - 2.4347e-1 (1.16e-1) $\approx$ 2.7857e-1 (1.26e-1) + 5.3432e-2 (3.94e-2) - 5.6590e-2 (4.69e-2) - 1.6657e-1 (1.25e-1)
5 10 1.1540e-1 (7.24e-2) - 2.2795e-1 (1.21e-1) $\approx$ 3.6663e-1 (1.53e-1) + 7.0922e-2 (4.65e-2) - 8.0312e-2 (4.85e-2) - 1.9786e-1 (1.18e-1)
10 10 7.4476e-1 (4.87e-2) - 7.6957e-1 (7.46e-2) - 8.6964e-1 (5.86e-2) + 6.2825e-1 (7.80e-2) - 7.4705e-1 (4.32e-2) - 8.3492e-1 (5.85e-2)
MaF6 3 10 2.8335e-4 (1.30e-3) - 1.0583e-1 (3.79e-2) - 1.6642e-1 (1.04e-2) + 3.1905e-3 (8.83e-3) - 1.0044e-2 (1.26e-2) - 1.3200e-1 (2.81e-2)
5 10 4.7388e-3 (1.19e-2) - 1.0321e-1 (1.15e-2) + 1.1474e-1 (2.86e-3) + 1.4998e-2 (1.57e-2) $\approx$ 2.6747e-2 (2.40e-2) $\approx$ 1.6809e-2 (2.57e-2)
10 10 9.6658e-2 (1.43e-3) - 9.8905e-2 (1.65e-3) - 9.8202e-2 (9.16e-4) - 9.3467e-2 (1.49e-3) - 9.7198e-2 (1.03e-3) - 9.9933e-2 (3.64e-4)
MaF7 3 10 2.5209e-1 (3.70e-3) - 2.5443e-1 (2.54e-2) - 2.5836e-1 (2.44e-2) - 2.6660e-1 (5.26e-3) - 2.6983e-1 (3.14e-3) - 2.8034e-1 (9.87e-4)
5 10 2.1118e-1 (6.01e-3) - 2.3450e-1 (1.16e-2) $\approx$ 2.3073e-1 (1.15e-2) $\approx$ 2.4569e-1 (6.20e-3) + 2.3236e-1 (6.93e-3) $\approx$ 2.3450e-1 (4.81e-3)
10 10 1.7515e-1 (5.57e-3) - 1.6996e-1 (1.20e-2) - 1.7435e-1 (6.23e-3) - 1.4747e-1 (8.54e-3) - 1.4603e-1 (7.26e-3) - 1.8201e-1 (8.98e-3)
MaF10 3 10 1.4257e-1 (4.47e-2) $\approx$ 1.5406e-1 (5.85e-2) $\approx$ 1.5921e-1 (4.38e-2) $\approx$ 2.9224e-2 (3.95e-2) - 2.0331e-2 (2.76e-2) - 1.3722e-1 (4.18e-2)
5 10 1.9544e-1 (3.39e-2) - 1.9753e-1 (6.91e-2) $\approx$ 1.9958e-1 (5.49e-2) $\approx$ 1.2859e-1 (5.06e-2) - 1.2046e-1 (3.56e-2) - 2.2253e-1 (5.30e-2)
10 10 2.2149e-1 (1.69e-2) - 6.4517e-1 (2.65e-1) - 6.2589e-1 (2.32e-1) - 2.1032e-1 (6.87e-2) - 2.2924e-1 (1.07e-2) - 8.7050e-1 (9.36e-2)
MaF11 3 10 7.4666e-1 (5.05e-2) - 7.4308e-1 (3.43e-2) - 8.2768e-1 (1.53e-2) - 7.9595e-1 (1.76e-2) - 8.0300e-1 (1.11e-2) - 8.5546e-1 (1.47e-2)
5 10 8.6388e-1 (4.10e-2) - 8.3081e-1 (3.64e-2) - 9.0993e-1 (4.04e-2) $\approx$ 9.0046e-1 (1.37e-2) - 9.0298e-1 (1.78e-2) - 9.2302e-1 (2.40e-2)
10 11 9.6832e-1 (1.09e-2) - 9.7324e-1 (1.05e-2) - 9.5402e-1 (4.88e-2) - 9.9164e-1 (2.64e-3) $\approx$ 9.7918e-1 (2.42e-2) - 9.8989e-1 (4.82e-3)
MaF12 3 10 2.3963e-1 (2.90e-2) - 2.8325e-1 (4.02e-2) $\approx$ 3.4154e-1 (3.96e-2) + 2.8119e-1 (2.98e-2) $\approx$ 2.9240e-1 (3.22e-2) $\approx$ 2.9569e-1 (3.25e-2)
5 10 3.5976e-1 (3.97e-2) - 4.3480e-1 (5.47e-2) $\approx$ 5.0032e-1 (5.80e-2) + 4.2081e-1 (5.94e-2) $\approx$ 4.1019e-1 (6.53e-2) $\approx$ 4.4101e-1 (6.63e-2)
10 10 6.5057e-1 (5.00e-2) - 7.4022e-1 (2.23e-2) $\approx$ 7.6448e-1 (1.90e-2) $\approx$ 5.9450e-1 (5.31e-2) - 6.2427e-1 (4.95e-2) - 7.4714e-1 (3.62e-2)
MaF13 3 10 1.6516e-1 (4.87e-2) - 3.6088e-1 (4.20e-2) - 3.9093e-1 (2.64e-2) $\approx$ 4.4328e-1 (1.29e-2) + 5.1229e-1 (4.68e-3) + 3.9014e-1 (2.20e-2)
5 10 0.0000e+0 (0.00e+0) - 9.8391e-2 (3.32e-2) + 6.0588e-2 (3.13e-2) + 1.0549e-1 (2.63e-2) + 5.2907e-4 (1.50e-3) - 1.7688e-2 (1.67e-2)
10 10 0.0000e+0 (0.00e+0) - 1.2057e-2 (1.61e-2) + 1.3949e-2 (1.61e-2) + 3.9530e-2 (2.86e-2) + 0.0000e+0 (0.00e+0) - 4.7121e-3 (8.03e-3)
+/-$\approx$ 0/27/6 4/16/13 9/11/13 6/18/9 3/22/8
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Table S-III: IGD+ RESULTS (MEAN AND STANDARD DEVIATION) OBTAINED BY THE PROPOSED CI-EMO AND THE VARIANT WITH SAME WEIGHTS